a football team conists of 20 each freshmen and sophmores, 15 juniors, and 10 seniors. four players are selcted at random to serve as captain find the probability that
a. all 4 are seniors
b. there is 1 each
c. there are 2 soph and 2 freshman
d. at least 1 is a senior

Answer by stanbon(75887)

a football team conists of 20 each of freshmen and sophmores, 15 juniors, and 10 seniors. four players are selcted at random to serve as captain find the probability that
a. all 4 are seniors
Ans: 10C4/65C4
----------------------------
b. there is 1 each
Ans: [20*20*15*10]/65C4 = 60,000/677040
----------------------------
c. there are 2 soph and 2 freshman
[20C2*20C2]/[65C4]
------------------------------
d. at least 1 is a senior
P(at least one senior) = 1-P(no senior) = 1 - [55C4/65C4]
============================
Cheers,
Stan H.
==========